Calculate the cash cycle based on the following information:

Sales = $146,625
COGS = $29,097
Inventory = $10,853
Accounts Receivable = $35,971
Accounts Payable = $10,456

Enter your answer rounded off to two decimal points.
Answer #1
Inventory turnover = COGS/inventory
Inventory turnover = 29097/10853
Inventory turnover = 2.68
days of inventory on hand = number of days in a year/inventory turnover
days of inventory on hand = 365/2.68
days of inventory on hand = 136.19
Receivables turnover = Credit sales/receivables
Receivables turnover = 146625/35971
Receivables turnover = 4.08
days of sales outstanding = number of days in a year/receivables turnover
days of sales outstanding = 365/4.08
days of sales outstanding = 89.46
Accounts payables turnover = purchases/payables
Accounts payables turnover = 29097/10456
Accounts payables turnover = 2.78
days of payables outstanding = number of days in a year/accounts payable turnover
days of payables outstanding = 365/2.78
days of payables outstanding = 131.29
Operating cycle = days of sales outstanding + days of inventory on hand
Operating cycle = 89.46+136.19
Operating cycle = 225.65
Cash conversion cycle = Operating cycle - days of payables outstanding
Cash cycle = 225.65-131.29
Cash cycle = 94.36